how do yu solve 6x+4y=-12 and 8x-14y=-16
Answer by checkley77(12844)

6x+4y=-12 multiply by 8.
8x-14y=-16 multiply by -6 & add.
48x+32y=-96
-48x+84y=96
-----------------
116y=0
y=0 ans.
6x+4*0=-12
6x=-12
x=-12/6
x=-2 ans.
Proof:
8*-2-14*0=-16
-16=-16